OpenStudy (anonymous):

Factor the expression completely: 4x^2 - 8x - 140 & x^2 - 64 Show work

OpenStudy (anonymous):

@e.mccormick @austinL @AkashdeepDeb @genius12 @myininaya Please help, I need this ASAP

OpenStudy (anonymous):

Given the function: f(x) = 4(x - 5)^2 + 6 First, identify the vertex: Then tell how each shift will affect the graph: f(x + 1) f(x) + 2 f(-3x)

OpenStudy (anonymous):

@tHe_FiZiCx99 @thomaster Please help

OpenStudy (akashdeepdeb):

First question 4x^2 - 8x - 140 4(x2 - 2x - 35) 4(x2 - 7x + 5x - 35) [SPLIT THE MIDDLE TERM] 4{x(x-7) + 5(x-7)} 4(x+5)(x-7) ULTIMATE FACTOR FORM! \m/

OpenStudy (akashdeepdeb):

For the second question use this theorem. \[a^2 - b^2 = (a+b)(a-b)\] The 2 terms are perfect squares so it should be easy. :)

OpenStudy (akashdeepdeb):

Third question f(x) = 4(x - 5)^2 + 6 Simplify the equation to get the individual a,b,c standard values. This is a parabola and thus to find the vertex you may use the vertex co-ordinate formula => V(x,y) = V(\(\frac{-b}{2a} , c - \frac{b^2}{4a}\))

OpenStudy (akashdeepdeb):

Fourth question Anything done to the x value will shift it in the opposite direction by the same number of units added/subtracted along the x-axis. Thus f(x+1) should move f(x) 1 unit to the left.

OpenStudy (akashdeepdeb):

Fifth question Anything added/subtracted to f(x) would move the entire graph, the same number units up or down along the y-axis. Thus f(x) +2 will move it 2 units upwards along the y-axis.

OpenStudy (akashdeepdeb):

Sixth question When anything is multiplied to the x value it makes it narrower, if it is divided by any magnitude it'll become more wider. When f(x) becomes f(-x) it mirrors itself along the y-axis. ie. -x^3 is the mirror image of x^3 and same with all other functions! So f(-3x) would be narrower and the mirror image of the original graph f(x)

OpenStudy (akashdeepdeb):

To test out all the answers, try it on, https://www.desmos.com/calculator Understood all of this? :)